Solve the equation of x²p²+xp-(y²+y)=0

1
Expert's answer
2022-02-08T14:47:57-0500

"\displaystyle
x²p²+xp-(y²+y)=0\\\Rightarrow p=\frac{-x\pm\sqrt{x^2-4x^2(-(y^2+y))}}{2x^2}=\frac{-x\pm\sqrt{x^2-4x^2(-y^2-y)}}{2x^2}"

"\displaystyle
=\frac{-x\pm\sqrt{x^2+4x^2(y^2+y)}}{2x^2}=\frac{-x\pm\sqrt{x^2(1+4(y^2+y))}}{2x^2}"

"\displaystyle
=\frac{-x\pm x\sqrt{(1+4(y^2+y))}}{2x^2}=\frac{-x\pm x\sqrt{4y^2+4y+1}}{2x^2}"

"\displaystyle
=\frac{-1\pm \sqrt{4y^2+4y+1}}{2x}", using quadratic formula.


"\displaystyle
\Rightarrow p=\frac{\partial z}{\partial x}=\frac{-1\pm \sqrt{4y^2+4y+1}}{2x}=\frac{-1\pm \sqrt{4y^2+4y+1}}{2x}=\frac{-1\pm\sqrt{(2y+1)^2}}{2x}"


"\displaystyle
\Rightarrow \frac{\partial z}{\partial x}=\frac{-1\pm(2y+1)}{2x}=
\begin{cases}
\frac{y}{x}& \text{if }\frac{\partial z}{\partial x}=\frac{-1+(2y+1)}{2x}\\
\quad\\
\frac{-1-y}{x}& \text{if }\frac{\partial z}{\partial x}=\frac{-1-(2y+1)}{2x}
\end{cases}"

"\displaystyle
\Rightarrow z=
\begin{cases}
\int\frac{y}{x}\ dx=y\int\frac{1}{x}\ dx=y\ln x+f(y)\\
\quad\\
\int\frac{-(1+y)}{x}\ dx=-(1+y)\int\frac{1}{x}\ dx=-(1+y)\ln x+ f(y)
\end{cases}"

where "f(y)" is an arbitrary function of "y".


Thus, "\displaystyle
z=y\ln x+f(y) \text{ or }z=-(1+y)\ln x+f(y)"